A 30.21 g sample of Ar (Argon) gas has a volume of 900 mL at a temperature of 200°C.
i. Calculate the pressure using the Ideal Gas Law.
ii. Calculate the corrected pressure using the Van der Waals Equation.
a=1.336 atm×L2 b=0.03201 L
mol2 mol
1
Expert's answer
2020-04-18T06:45:14-0400

Molar mass of Argon is 40 gms.

Thus, number of moles "= n=30.21\/40=0.755"


Given : "V=0.9 L ; T=200\u00b0C=273+200=473K"

We know Gas constant "= R=0.0821 Latm mol^{-1}K^{-1}"


i. Ideal gas law : "PV=nRT \implies P=nRT/V"

"P=0.755*0.0821*473\/0.9=32.577 atm"---(Answer)


ii. Van der Waals Equation : "(P+an^2\/V^2)(V-nb)=nRT"

Given : "a=1.336atmL^2 mol^{-2}; b=0.03201L mol^{-1}"

Thus, "(P+(1.336*0.755^2\/0.9^2))(0.9\u2212(0.755*0.03201)=0.755*0.0821*473"

"\implies (P+0.94)(0.9-0.723)=29.319"

"\implies P=(29.319/0.177)-0.94=164.704 atm" ---(Answer)